How do I work this problem -6+4/3(2)-1= I tried working it by first doin the multiplication of 3(2) and simplifing 4/6 to 2/3, next I tried adding -6+-1 and got -7. That's as far as I got.

Answer by rfer(16322)

multiply the complete fraction
-6+4/3(2)-1
-6+8/3-1
-18/3+8/3-3/3=-13/3=-4 1/3